A ball is rolling down a hill with an initial velocity of 9m/s. What is the acceleration of the ball if its displacement is 10m in 4s?
1
Expert's answer
2020-11-05T10:44:34-0500

The displacement of a body moving with a constant acceleration is:


"d = v_0t + \dfrac{at^2}{2}"

where "v_0 = 9m\/s" is the initial velocity, "t = 4s" is the time of movment and "a" is the acceleration. Expressing "a", obtain:


"a = \dfrac{2d}{t^2} - \dfrac{2v_0}{t} = \dfrac{2\cdot 10m}{(4s)^2} - \dfrac{2\cdot 9m/s}{4s} = -3.25m/s^2"

Answer. -3.25 m/s^2